OpenStudy (anonymous):

12½ % of a number P is 70 Find the number P

OpenStudy (anonymous):

12 1/2%=12.5% 12.5% x p = 70 p=70 /12.5% p=70/ (12.5/100) p=70 x 100 / 12.5 p=560

OpenStudy (anonymous):

Wow!! Thanks very much!! GOD BLESS YOU!